Please help me solve these problems. Solve the systems of linear equations.
1/2x + 1/3y= 1/3 1/6x+1/3y=2/9

4x +10y=-28 6x+7y=6

-1.25x+0.5y=-1.6875 0.2x-0.5y=0.9
solve the equation
4x-3y=4 12x-9y=12

Answer by Alan3354(69443)

Please help me solve these problems. Solve the systems of linear equations.
1/2x + 1/3y= 1/3
1/6x+1/3y=2/9
-------
Multiply by whatever it takes to eliminate the fractions.
-------------
4x +10y=-28
6x + 7y =6
Don't put 2 equations on one line. Use something to separate problems.
================================
-1.25x+0.5y=-1.6875 0.2x-0.5y=0.9
solve the equation
4x-3y=4
12x-9y=12
Divide the 2nd eqn by 3, compare it to the 1st.